# Partonic distribution functions and amplitudes using tensor network methods

### v2

@@ 3 Results
In this work, we prepare the hadron state and simulate the real-time correlators of the quark and gluon fields using tensor network framework ITensor ITensor:2022 . The numerical results are compared with available exact diagonalization and quantum simulation results. Throughout this paper, we always use the lattice spacing $a=1$ in our simulations. Importantly, our simulation follows the implementation using DMRG for state preparation and the TDVP for time evolution; whereas the recent work on Schwinger model uses DMRG and TEBD respectively Banuls:2024oxa .
@@ 3.1 Hadron state preparation
Here, within the 1+1 dimension NJL model, our hadronic states are prepared via the DMRG algorithm. To prepare the hadron state on a tensor network within the neutral charge $Q=0$ sector, mimicking a charge-neutral pion ($\pi^{0}$), we utilize a charge-preserving (CP) lattice. With our setup, the total charge operator is defined as $Q=\sum_{i=1}^{N}(\sigma^{z}_{n}+(-1)^{n})/2$ on a lattice of $N$ qubits Ikeda:2024rzv . It is important to note that when working on a CP lattice, the Hamiltonian operator should be written in a form where its terms consist of only $\sigma^{z}$ or both $\sigma^{-},\sigma^{+}$ to ensure charge conservation, as in eq. (13). We first obtain the vacuum state (i.e. the ground state) on a CP lattice with sufficient number of sweeps and numerical accuracy. Specifically, we found 100 sweeps and 200 bond dimensions sufficient for simulating a system of 100 qubits. The high sweeps are found to be crucial to ensure DMRG convergence for simulation at larger qubits.
To obtain the hadron state (i.e., the first excited state), we project out the vacuum MPS and perform another set of DMRG sweeps with sufficiently large weights. For the hadron state, we found that even higher sweeps and maximum bond dimension are necessary to reach convergent mass results. Specifically, we used 500 sweeps and 800 bond dimensions. Throughout our simulation, we ensured that the obtained MPS consistently maintained a total charge of zero and that the bond dimension remained within the allocated maximum limit. In Table. 1, we summarize our hadron mass $M_{h}$, defined as the mass difference between the excited state and the ground state, findings for various system sizes $N$:
$M_{h}=\langle h|H|h\rangle-\langle\Omega|H|\Omega\rangle\,,$ (17)
Specifically, each value of $N$ is picked such that the two boundaries are matched up, i.e., $\bar{N}_{\mathrm{max}}=-\bar{N}_{\mathrm{min}}$. As the system size increases, $M_{h}$ becomes stable and the DRMG algorithm is able to identify the mass gap with high accuracy up to $10^{-8}$. For this reason, we always use $N=102$ qubits in preparing the ground and excited states for the final results.
Table 1: Hadron masses obtained from DMRG on CP lattice for different qubits. Here, $m=0.7a^{-1}$ and $g=0.25$. The DMRG results are accurate up to $10^{-8}$ with sufficient sweeps (100-500) and bond dimensions (200-800) at each qubit size.
@@ 3.2 Real-time evolution
where |$\ket{\psi_{L}},\ket{\psi_{R}}$ can be either |$\ket{h}$ or |$\ket{\Omega}$. Since |$\ket{\psi_{L}}$ is an eigenstate of $H$ (with eigenvalue $E_{L}$), the left-side time evolution operator reduces to an overall phase factor that greatly simplifies the computation [^3] We verified numerically the two expressions for $E_{ab}(z)$ in eq. (18) yield identical results.. Applying the local operators on MPS is straightforward and directly built into ITensor that involves three designated procedures: contraction, unpriming, and insertion of the MPS tensor. To time evolve the MPS, we use the TDVP algorithm Haegeman:2011zz ; haegeman2016unifying , included with the ITensorTDVP library. Specifically, we always use one time sweep and a maximum bond dimension of 810 (slightly larger than our max hadron state bond dimension). We tested the effect of different number of TDVP sweeps on the simulation results and found no significant improvement with more than one sweep. For large system sizes, we note that both the bond dimension of the MPS state and cost of the simulation increase at longer evolution time $z$. To avoid hitting the allocated maximum bond dimension and reduce computational complexity, we simulate in the region between $z=-10$ and $z=10$ where the matrix elements have substantial non-zero contributions. We verified that simulating beyond this region yielded no significant improvement to our final results.
Besides using the TDVP algorithm, we also implemented time evolution using the TEBD approach. At its heart, TEBD implements the Suzuki trotterization with approximation in the MPS. In comparison with the TDVP, we found it to be less accurate and more computationally demanding.
@@ 3.3 Parton distribution function and amplitudes
For small system sizes, we find consistent agreement among the TN approach, exact diagonalization (ED), and quantum circuit (QC) simulation. For the quantum circuit simulation, we used first-order trotterization Trotter1959 to implement time evolution with scaling $O(\tau^{2})$ for the time step $\tau$ that matches the order of TEBD and TDVP childs2019nearly ; childs2021theory . State preparations for the QC simulation is dramatically simplified and taken as exact eigenstates. In Fig. 1, we present the numerical results for the PDF using a lattice of $N=10$ qubits. From Fig. 1(a), we see a strong alignment among all three results, implying that our setup is valid and generally applicable to quantum computing. In Fig. 1(b), we make a log-linear plot of the absolute error $\Delta f_{q}(x)$ in TN and QC methods compared with exact diagonalization for each $x$. Noticeably, the TN consistently and significantly reduces the error by six orders of magnitude with an average error of around $10^{-8}$, which aligns with the numerical accuracy of the DMRG method for state preparation. Similar observations are found in computing the DA at $N=10$ qubits. These observations prove the reliability of the TN strategies used in our numerical calculations.
@@
To demonstrate numerical results in the continuum limit, we use sufficiently high number of qubits in our simulations to compute the PDF and DA using tensor networks. Such calculation is beyond the current limits of exact simulation and quantum simulation with classical hardware. With tensor networks, we are able to study physical results at continuum limits utilizing an efficient representation of the many-body system with the MPS. In Fig. 2, we compute both the PDF and DA in the same setting with increasing system sizes (i.e., qubits) using selected mass $m=0.7a^{-1}$ and strong coupling $g=0.25$. For the PDFs, as expected in a one-flavor hadron (such as $\pi^{0}$, composed of many pairs of $q$ and $\bar{q}$), we see an asymmetric plot with a peak at $x=0.5$ and $x=-0.5$. Similarly for the DAs, we see a typical distribution centered in the $x\in[0,1]$ region. As system size $N$ increases, both the PDF and DA results gradually converge to the large-qubit curve at $N=102$, suggesting the calculations approach the continuum of $N\rightarrow\infty$ limit [^4] In principle, one would take both $a\rightarrow 0$ and $N\rightarrow\infty$. For simplicity, we fix $a=1$.. Therefore, for all the results below, we always use $N=102$ qubits for the simulation and we find it sufficient to represent the continuum behavior of the PDF and DA at this large qubit limit for our NJL Hamiltonian. Overall, the shape and tendency of our results are qualitatively consistent with experimental observations such as E791 data E791:2000xcx for the PDF and Fermilab E615 data Conway:1989fs for the DA of the pion and they also resemble numerical calculations in 1+1 QCD Jia:2018qee .
@@
Figure 2: Simulation results for the PDF and DA over increasing number of qubits $N$. Here, we fix $m=0.7a^{-1}$ and $g=0.25$. The asymptotic of perturbative QCD for the DA is provided.
One important strength of our approach is the freedom to investigate the PDF and DA at various coupling strengths, especially in the non-perturbative regions that are generally complicated. In Fig. 3, we show the PDF and DA results at various coupling strengths $g=0.25,0.5,1.0$ with the same $m=0.5a^{-1}$ and $N=102$ qubits. Interestingly, the PDFs show distinctive behavior at different coupling strengths. At small coupling $g=0.25$, we get a sharp peak at $x=0.5$ and $x=-0.5$ that is consistent with the perturbative QCD expectation. With increased coupling strength, the PDF becomes less sharply peaked. Similarly for the DA, at the low coupling such as $g=0.25$, the distribution is sharply centered at $x=0.5$, comparable to the perturbative limit of $\phi_{\mathrm{asym}}(x)=6x(1-x)$ Lepage:1980fj . When we turn on the coupling strength towards $1$, we see significant non-perturbative effects, giving broadened DA curves. Qualitatively, our results are also in agreement with available quantum and tensor network simulation results from the light-front Hamiltonian methods Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp and lattice calculations LiTianyin:2021kcs ; LiTianyin:2024nod ; Banuls:2024oxa .
@@ 4 Conclusion and outlook
In this work, we used a quantum strategy based on tensor network simulation to extract the PDF and DA for hadrons in the Nambu-Jona-Lasinio model in 1+1 dimensions in the same setting. Our approach utilizes the DMRG algorithm to accurately construct the hadron and vacuum states on a charge-preserving lattice and the TDVP algorithm to efficiently time evolve the two-point correlator needed for these non-perturbative quantities. With our approach, we extract the PDF and DA in the same setting with converging results up to 102 qubits. We are able to study the behavior of the PDF and DA under the same setting for various masses and strong couplings, and observe consistent agreement with the expectations from the perturbative QCD as well as non-relativistic limit. Our exploratory study shows the advantage of tensor network in the current era for computing various physically important quantities in the continuum limit.
It would be very important to expand non-perturbative calculations of quantities such as the PDF, DA, and hadronic tensors using tensor network to U(1) and SU(2) gauge theories. Notably, similar calculation using TN methods has been successfully demonstrated in Ref. Banuls:2024oxa for the U(1) case in the Schwinger model. Inclusion of multiple flavor and spin degrees of freedom in the line of study would provide significant understanding to realistic mesons and baryons from first principles. While tensor networks have primarily been applied to 1+1 dimensions, it would be interesting to extend to higher lattice dimensions. The natural extension of MPS to higher dimensions would be projected entanglement pair states Verstraete:2004cf ; verstraete2004valence , tree tensor network Shi:2005ctp , and so forth, which have been used to study 3+1 dimension Lattice Gauge Theory Magnifico:2020bqt ; Banuls:2018jag . We also plan to incorporate more advanced simulation techniques, such as tensor renormalization group LevinNave:2007 and its natural extension, tensor network renormalization tnr , to scale to higher qubit simulations with reduced error and study entanglement structure. These methods can be used to calculate correlation functions and generalize to systems beyond one spatial dimension.

### v3

@@ 3 Results
In this work, we prepare the hadron state and simulate the real-time correlators of the quark and gluon fields using tensor network framework ITensor ITensor:2022 . The numerical results are compared with available exact diagonalization and quantum simulation results. Throughout this paper, we always use the lattice spacing $a=1$ in our simulations. Importantly, our simulation follows the implementation using variational ground-state search algorithm for state preparation and the TDVP for time evolution; whereas the recent work on Schwinger model uses TEBD for the time evolution Banuls:2025wiq .
@@ 3.1 Hadron state preparation
Here, within the 1+1 dimension NJL model, our hadronic states are prepared via two-site varitional ground-state search algorithm. To prepare the hadron state on a tensor network within the neutral charge $Q=0$ sector, mimicking a charge-neutral pion ($\pi^{0}$), we utilize a charge-preserving (CP) lattice to conserve the quantum number. With our setup, the total charge operator is defined as $Q=\sum_{i=1}^{N}(\sigma^{z}_{n}+(-1)^{n})/2$ on a lattice of $N$ qubits Ikeda:2024rzv . When applying variational algorithms to obtain both the vacuum state and hadron state, we start with a guess ansatz that we know analytically has $Q=0$; then the ground-state search algorithm ensures that we remain in the $Q=0$ sector. It is important to note that when working on a CP lattice, the Hamiltonian operator should be written in a form where its terms consist of only $\sigma^{z}$ or both $\sigma^{-},\sigma^{+}$ to ensure charge conservation, as in eq. (13). We first obtain the vacuum state (i.e. the ground state) on a CP lattice with sufficient number of sweeps and numerical accuracy. Specifically, we found 100 sweeps and 200 maximum bond dimensions sufficient for simulating a system of 100 qubits at a cutoff of $10^{-12}$. By sufficient we mean that the bond dimension was not fully exhausted during the variational search. The high sweeps are found to be crucial to ensure numerical convergence for simulation at larger qubits.
To obtain the hadron state (i.e., the first excited state), we project out the vacuum MPS and perform another set of sweeps with sufficiently large weights. For the hadron state, we found that even higher sweeps and maximum bond dimension are necessary to reach convergent mass results. Specifically, we used up to 500 sweeps and 800 maximum bond dimension. Throughout our simulation, we ensured that the obtained MPS consistently maintained a total charge of zero and that the bond dimension remained well below the allocated maximum limit. In Table. 1, we summarize our hadron mass $M_{h}$, defined as the mass difference between the excited state and the ground state, findings for various system sizes $N$:
$M_{h}=\langle h|H|h\rangle-\langle\Omega|H|\Omega\rangle\,.$ (18)
Specifically, each value of $N$ is picked such that the two boundaries are matched up, i.e., $\bar{N}_{\mathrm{max}}=-\bar{N}_{\mathrm{min}}$. As the system size increases, $M_{h}$ becomes stable and we are able to identify the mass gap with high accuracy. For this reason, we always use $N=102$ qubits in preparing the ground and excited states for the final results. Note the hadron mass values in Table. 1 are converged results obtained with sufficient maximum bond dimension with a cutoff of $10^{-12}$. See Table. 2 in Appendix 1 for a detailed summary of the hadron mass values at various numerical cutoffs and maximum bond dimensions that shows our results converge with increasing maximum bond dimensions and cutoffs, respectively.
Table 1: Hadron masses obtained from variational ground-state search algorithm on CP lattice for different qubits. Here, $m=0.7a^{-1}$ and $g=0.25$. The results are accurate and converged with sufficient sweeps (100-500) and maximum bond dimensions (200-800) at a cutoff $10^{-12}$ for each qubit size.
@@ 3.3 Parton distribution function and amplitudes
For small system sizes, we find consistent agreement among the TN approach, exact diagonalization (ED), and quantum circuit (QC) simulation. For the quantum circuit simulation, we used first-order trotterization Trotter1959 to implement time evolution with scaling $\mathcal{O}(\tau^{2})$ for the time step $\tau$ that matches the order of TDVP childs2019nearly ; childs2021theory . Specifically, we use the statevector simulator from Qiskit Qiskit library to perform exact simulation with no sampling noise from shots. State preparations for the QC simulation is dramatically simplified and taken as exact eigenstates directly solved from diagonalization. In Fig. 1, we present the numerical results for the PDF using a lattice of $N=10$ qubits. From Fig. 1(a), we see a strong alignment among all three results, implying that our setup is valid and generally applicable to quantum computing. In Fig. 1(b), we make a log-linear plot of the absolute error $\Delta f_{q}(x)$ in TN and QC methods compared with exact diagonalization for each $x$. Noticeably, the TN consistently and significantly reduces the error by six orders of magnitude with an average error of around $10^{-8}$, which aligns with the numerical accuracy of the variational method for state preparation. Similar observations are found in computing the DA at $N=10$ qubits. These observations prove the reliability of the TN strategies used in our numerical calculations.
@@
To demonstrate numerical results in the continuum limit, we use sufficiently high number of qubits in our simulations to compute the PDF and DA using tensor networks. Such calculation is beyond the current limits of exact simulation and quantum simulation with classical hardware. With tensor networks, we are able to study physical results at continuum limits utilizing an efficient representation of the many-body system with the MPS. In Fig. 2, we compute both the PDF and DA in the same setting with increasing system sizes (i.e., qubits) using selected mass $m=0.7a^{-1}$ and strong coupling $g=0.25$. For the PDFs, as expected in a one-flavor hadron (such as $\pi^{0}$, composed of many pairs of $q$ and $\bar{q}$), we see an asymmetric plot with a peak at $x=0.5$ and x=0.5x=-0.5. Similarly for the DAs, we see a typical distribution centered in the $x\in[0,1]$ region. As system size $N$ increases, both the PDF and DA results gradually converge to the large-qubit curve at $N=102$, suggesting the calculations approach the continuum of $N\rightarrow\infty$ limit [^4] In principle, one would take both $a\rightarrow 0$ and $N\rightarrow\infty$. For simplicity, we fix $a=1$.. Therefore, for all the results below, we always use $N=102$ qubits for the simulation and we find it sufficient to represent the continuum behavior of the PDF and DA at this large qubit limit for our NJL Hamiltonian. Overall, the shape and tendency of our results are qualitatively consistent with experimental observations such as E791 data E791:2000xcx for the PDF and Fermilab E615 data Conway:1989fs for the DA of the pion and they also resemble numerical calculations in 1+1 QCD Jia:2018qee .
@@
Figure 2: Simulation results for the PDF and DA over increasing number of qubits NN. Here, we fix $m=0.7a^{-1}$ and $g=0.25$. The asymptotic of perturbative QCD for the DA is provided.
One important strength of our approach is the freedom to investigate the PDF and DA at various coupling strengths, especially in the non-perturbative regions that are generally complicated. In Fig. 3, we show the PDF and DA results at various coupling strengths $g=0.25,0.5,1.0$ with the same $m=0.5a^{-1}$ and $N=102$ qubits. Interestingly, the PDFs show distinctive behavior at different coupling strengths. At small coupling $g=0.25$, we get a sharp peak at $x=0.5$ and x=0.5x=-0.5 that is consistent with the perturbative QCD expectation. With increased coupling strength, the PDF becomes less sharply peaked. Similarly for the DA, at the low coupling such as $g=0.25$, the distribution is sharply centered at $x=0.5$, comparable to the perturbative limit of $\phi_{\mathrm{asym}}(x)=6x(1-x)$ Lepage:1980fj . When we turn on the coupling strength towards $1$, we see significant non-perturbative effects, giving broadened DA curves. Qualitatively, our results are also in agreement with available quantum and tensor network simulation results from the light-front Hamiltonian methods Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp and lattice calculations LiTianyin:2021kcs ; LiTianyin:2024nod ; Banuls:2024oxa ; Banuls:2025wiq .
@@ 4 Conclusion and outlook
In this work, we used a quantum strategy based on tensor network simulation to extract the PDF and DA for hadrons in the Nambu–Jona-Lasinio model in 1+1 dimensions in the same setting. Our approach utilizes the variational ground-state search algorithm to accurately construct the hadron and vacuum states on a charge-preserving lattice and the TDVP algorithm to efficiently time evolve the two-point correlator needed for these non-perturbative quantities. With our approach, we extract the PDF and DA in the same setting with converging results up to 102 qubits. We are able to study the behavior of the PDF and DA under the same setting for various masses and strong couplings, and observe consistent agreement with the expectations from the perturbative QCD as well as non-relativistic limit. Our exploratory study shows the advantage of tensor network in the current era for computing various physically important quantities in the continuum limit.
It would be very important to expand non-perturbative calculations of quantities such as the PDF, DA, and hadronic tensors using tensor network to U(1) and SU(2) gauge theories. Notably, similar calculation using TN methods has been successfully demonstrated in Ref. Banuls:2025wiq for the U(1) case in the Schwinger model. Inclusion of multiple flavor and spin degrees of freedom in the line of study would provide significant understanding to realistic mesons and baryons from first principles. While tensor networks have primarily been applied to 1+1 dimensions, it would be interesting to extend to higher lattice dimensions. The natural extension of MPS to higher dimensions would be projected entanglement pair states Verstraete:2004cf ; verstraete2004valence , tree tensor network Shi:2005ctp , and so forth, which have been used to study 3+1 dimension Lattice Gauge Theory Magnifico:2020bqt ; Banuls:2018jag . We also plan to incorporate more advanced simulation techniques, such as tensor renormalization group LevinNave:2007 and its natural extension, tensor network renormalization tnr , to scale to higher qubit simulations with reduced error and study entanglement structure. These methods can be used to calculate correlation functions and generalize to systems beyond one spatial dimension.
@@ Acknowledgement
Appendix A Hadron mass details obtained from tensor network simulations
Table 2: Hadron masses obtained from variational ground-state search algorithm on CP lattice for different qubits with various truncation error cutoffs $\Delta$ and maximum bond dimensions $D_{\mathrm{max}}$. Here, $m=0.7a^{-1}$ and $g=0.25$ are used. For varying cutoffs, we use fixed $D_{\mathrm{max}}=800$; for varying maximum bond dimensions, we use fixed $\Delta=10^{-12}$. All simulations are done with sufficient number of sweeps (100-500) depending on system sizes.
Qubits (N) $18$ $30$ $50$ $102$
$\Delta=10^{-6}$ 1.64599824 1.64356479 1.64242277 1.64187559
$\Delta=10^{-8}$ 1.64600860 1.64359023 1.64246818 1.64188719
$\Delta=10^{-10}$ 1.64600864 1.64359027 1.64246829 1.64188740
$\Delta=10^{-12}$ 1.64600864 1.64359027 1.64248883 1.64188740
Qubits (N) $18$ $30$ $50$ $102$
$D_{\mathrm{max}}=40$ 1.64600864 1.64359029 1.64246831 1.64188746
$D_{\mathrm{max}}=80$ 1.64600864 1.64359027 1.64248690 1.64188740
$D_{\mathrm{max}}=100$ 1.64600864 1.64359027 1.64248918 1.64188740
$D_{\mathrm{max}}=200$ 1.64600864 1.64359027 1.64248883 1.64188740
$D_{\mathrm{max}}=400$ 1.64600864 1.64359027 1.64248883 1.64188740
$D_{\mathrm{max}}=800$ 1.64600864 1.64359027 1.64248883 1.64188740